# ZMP-HH/24-01 The Standard Model from String Theory: What Have We Learned?

### 03 Jan 2024 v1

@@ 4.3 MSSM and Supersymmetry Breaking
A classical scheme is the one found in heterotic compactifications, where one of the factors hosts the SM sector and the other one is a hidden sector. A gaugino condensate in the latter will generate a scalar potential that will both stabilize moduli and break supersymmetry, which is then communicated to the visible sector via gravitational effects Derendinger:1985kk (112, 113). The problem with this mechanism is that the generated potential has runaway directions toward a non-interacting theory Dine:1985he (114) and so extra ingredients must be added to generate a more involved potential with actual minima. One possibility is to add background fluxes, but their quantization implies that supersymmetry is broken at a high scale. Alternatively, one may consider several gauge factors that yield a gaugino condensate, so that the generated superpotential involves a sum of competing exponentials, also known as a racetrack scenario Krasnikov:1987jj (115). This idea can be made precise in toroidal orbifold settings, since there one can use modular invariance to compute the one-loop corrections to the gauge kinetic function, and from there the exact superpotential Font:1990nt (116, 117, 118). The racetrack scenario has however the disadvantage of typically leading to vacua with negative cosmological constant.
Type II orientifold compactifications provide a different scheme. Particularly interesting are Type IIB compactifications with O3/O7-planes. Here we have two different mechanisms that generate a scalar potential, and each of them acts on a different set of light fields, with no kinetic mixing at tree-level between both sets. The first mechanism consists of adding background three-form fluxes, which fix complex structure moduli at a high scale . The second one is the presence of D-brane instantons or perturbative corrections, which stabilize Kähler moduli at a lower scale. This structure has led to several proposals to build de Sitter vacua in string theory Kachru:2003aw (106, 107), whose implementation is currently under debate. One interesting property of this scheme is that three-form fluxes can in principle break supersymmetry at a scale much lower than , and that one can compute microscopically the effect of this supersymmetry breaking on the SM sector of the compactification, which is realized on the worldvolume of D-branes Camara:2003ku (119, 120, 121). This implies that, in Type II and F-theory constructions where the MSSM or extensions are localized on a patch of the extra dimensions, one can provide a geometric description of their soft term structure.
@@
From the macroscopic viewpoint, both the heterotic and Type II models fall into the category of gravity-mediated supersymmetry breaking. This is a general 4d supergravity framework in which the MSSM soft terms are parametrized by F-terms for bulk fields, which have Planck-suppressed couplings to the SM fields Kaplunovsky:1993rd (123, 124). In the simplest cases, realized by dilaton domination in the heterotic case and by -modulus domination in intersecting D7-branes/F-theory models, all soft terms are specified by only two parameters: the universal gaugino mass and the Higgs -term. This simple pattern has the advantage that it can easily avoid unwanted flavor-changing neutral currents, because the flavor patterns in the holomorphic Yukawa sector depend on a set of moduli that decouple from the supersymmetry-breaking sector Conlon:2007dw (125). Imposing experimental constraints on these models (like compatibility with EW symmetry breaking) reduces the two parameters to one, and leads to a scenario where the eventual measurement of a single sparticle mass determines the whole spectrum Aparicio:2008wh (126). In other constructions, like -modulus domination in heterotic and D3-branes at singularities models, the leading-order soft terms vanish, and they only appear through one-loop corrections to the gauge kinetic functions and Kähler potential. Such corrections are however quite model dependent, so it is difficult to draw general lessons.
Alternatives to gravity mediated supersymmetry breaking have also been explored in string theory setups, but their full realization remains an open problem. On the one hand, one-loop anomaly mediation requires to implement the idea of sequestering in the SM sector, which has been explored in Type IIB orientifolds with the SM localized on a warped throat Kachru:2007xp (127, 128). On the other hand, gauge mediation needs to localize the SM in a small region of the extra dimensions and full stabilization of the bulk fields of the compactification at a high scale Diaconescu:2005pc (129, 130, 131). These requirements are in tension with the realizations of the SM known to date.
@@ 4.4 Dark Sectors
One might wonder whether the appearance of dark sectors is a moduli dependent statement. If the hidden gauge sector contains massless charged scalar fields, its gauge algebra can in principle be partially higgsed, thereby reducing also the amount of matter, possibly even completely. However, so-called non-Higgsable clusters Morrison:2012np (134, 135), i.e. gauge sectors with no massless charged scalars that could induce a (partial) Higgsing, seem to be another generic property of the non-perturbative landscape of string vacua. This phenomenon can be studied explicitly in F-theory compactifications: An overwhelming prevalence of non-Higgsable clusters has been observed within a large set of elliptic Calabi-Yau fourfolds Halverson:2017ffz (136, 137). The glueballs of such confining dark sectors may give rise to interesting dark matter candidates though a typical challenge in this scenario is to prevent the potential oversaturation of the relic abundance. The phenomenological aspects of this scenario have been studied, for instance, in Halverson:2018olu (138, 139) and references therein.
The rich sequestered gauge sectors described above arise in addition to the matter sector associated with the higher-dimensional gravitational dynamics itself. Apart from the moduli and axion sector this may potentially even include the Kaluza-Klein towers of light fields such as the graviton. The idea of a light Klauza-Klein tower as the origin of dark matter was recently investigated in Gonzalo:2022jac (140) in the context of the Swampland approach to quantum gravity. It rests on the assumption that the smallness of the observed value of the cosmological constant points to our vacuum as arising in the asymptotic regions of a moduli space with one large extra dimension Montero:2022prj (141). An alternative, and possibly dual, dark matter candidate in such scenarios, and in string theory more generally, are primordial black holes, see Anchordoqui:2022txe (142) and references therein.
@@ 5 THE STANDARD MODEL AND QUANTUM GRAVITY
Interestingly, a different notion of naturalness has emerged from studying the vast but finite landscape of string vacua. Stringy naturalness Douglas:2004zg (148) is a measure of the degree of tuning by the number of string theory solutions leading to a given value of an observable. Naturalness based on the statistics of string vacua has been used to assess the likelihood of the scale of supersymmetry breaking Susskind:2004uv (149, 150, 151) and in addressing other vexing hierarchy problems such as the smallness of the cosmological constant, though explicit string theory realizations of this idea remain to be found.
@@
It was in this backdrop of finding how gravity affects quantum field theory reasonings that the Swampland Program was proposed, named after the term introduced in Vafa:2005ui (152). The goal of the Swampland Program is to discriminate between EFTs that belong to the Landscape (defined to be the set of theories that can descend from a quantum theory of gravity) and those that do not. See e.g. Brennan:2017rbf (153, 154, 155, 156, 157) for recent reviews. A number of quantum gravity constraints (also known as Swampland criteria) have been put forth, mostly by finding patterns in string constructions or through thought experiments involving black holes. The motive behind finding these constraints is to discover the underlying universal principles for consistently coupling quantum field theories to gravity. These quantum gravity constraints, if proven, can have interesting phenomenological consequences. For example, milli-charged dark matter scenarios frequently considered in phenomenological model building are in tension with the absence of global symmetry in quantum gravity Shiu:2013wxa (158). The Weak Gravity Conjecture (WGC) Arkani-Hamed:2006emk (159), which quantifies how much does gravity violate global symmetries, has been used to put phenomenological constraints on axions Brown:2015iha (160, 161, 162, 163, 96) and dark photons Reece:2018zvv (164). There have been attempts to leverage the WGC in the presence of scalars to set an upper bound on the weak scale Lust:2017wrl (165, 166). A stronger versions of the WGC which presumes the absence of stable non-supersymmetric AdS vacua Ooguri:2016pdq (167) has been used to link the observed value of the cosmological constant with the neutrino masses (which for fixed Yukawa couplings, set the weak scale) Ooguri:2016pdq (167, 168, 169). It should be noted that such arguments hinge on being able to demonstrate that the AdS vacua in question are stable. Without knowledge of the full UV spectrum, we cannot exclude the possibility of lower energy solutions. In fact, the Standard Model itself already contains a source of instability. Charged matter like the electron generates a potential for the Wilson line moduli (originated from the of the Standard Model upon compactification). It was pointed out in Hamada:2017yji (169) that the potential after taking into account the Wilson line moduli has a runaway behavior. Gravitational positivity bounds Hamada:2018dde (170, 171, 172, 173) obtained through studies of scattering amplitudes have recently been used to constrain new physics models involving massive gauge bosons Aoki:2023khq (174). Since the focus of this review is to extract lessons we learned from top-down constructions rather than bottom-up Swampland considerations, we refer the readers to some recent reviews Cvetic:2022fnv (4, 175, 176) for an in-depth discussion of other phenomenological consequences of the Swampland idea.
@@
As we have seen in the previous sections, the gauge and matter content of top-down string constructions are tightly constrained. An interesting question is whether the constrained spectra are artifacts of looking under the lamppost, or there are general principles that forbid certain particles or interactions to arise in a consistent quantum theory of gravity. Recent works have gathered support for the latter. The Completeness Hypothesis Polchinski:2003bq (177) asserts that physical states with all possible gauge charges consistent with Dirac quantization are present in a consistent theory of quantum gravity. It has since been found that the Completeness Hypothesis follows from the absence of generalized non-invertible global symmetries Gaiotto:2014kfa (178, 145, 179, 180). A corollary of this hypothesis when applied to states charged under 1-form gauge fields is the existence of magnetic monopoles. When higher-form symmetries come into play, the physical states in question would include extended objects, such as strings and branes. A novel approach using brane probes to test the consistency of EFTs coupled to gravity was initiated in Kim:2019vuc (181) where it was shown that unitarity on the brane probes can rule out infinite families of anomaly free gravitational theories. Further studies have found precise match of the allowed spectra with string constructions, sometimes involving the global structure of the spacetime gauge group Cvetic:2020kuw (182, 183, 184, 185, 186, 187, 188, 189). Taken together, these findings suggest a notion of string universality Adams:2010zy (190) (or the string lamppost principle Montero:2020icj (183)) that all consistent theories of quantum gravity are realized in string theory. It is conceivable that exploring this avenue of research to gravitational theories in lower dimensions and with reduced supersymmetry may shed light on the constrained spectra observed in four-dimensional string vacua that bear significance in phenomenological contexts. Before then, six-dimensional supergravity theories and their string/F-theory realizations are an intermediate testbed for substantiating this approach. This is because this class of string constructions is fairly well understood, and known quantum consistency conditions already tightly bound the set of possible low-energy theories. See Morrison:2021wuv (191, 192) for investigations of the Completeness Hypothesis and related quantum gravity constraints in this controlled setting.
@@
We hope we have conveyed that further understanding of quantum gravity constraints, in tandem with continued advances in string compactifications, has the potential to unravel the inherent structure found in the particle spectra prevalent in string theory.
@@ 6 WHAT HAVE WE LEARNED?
Quantum gravity is essential to understand particle physics: Gravitational physics plays a more important role for particle model building than naively expected, and string theory, unlike purely field theoretic approaches to phenomenology, has all the ingredients to help us understand this profound connection. For instance, all aspects of abelian or discrete gauge symmetries and their associated selection rules are sensitive to the gravitational sector and cannot be studied in isolation. More generally this is a key lesson from the Swampland Program. Harvesting the constraints from quantum gravity for particle physics via string theory as a computationally well defined framework furnishes a promising research direction for modern string phenomenology.
@@ ACKNOWLEDGMENTS
F. M. is supported through the grants CEX2020-001007-S and PID2021-123017NB-I00, funded by MCIN/AEI/10.13039/501100011033 and by ERDF A way of making Europe. G. S. is supported in part by the DOE grant DE-SC0017647. T. W. is supported in part by Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy EXC 2121 Quantum Universe 390833306 and by Deutsche Forschungsgemeinschaft through a German-Israeli Project Cooperation (DIP) grant “Holography and the Swampland”.

### 08 Feb 2024 v2

@@ 4.3 MSSM and Supersymmetry Breaking
A classical scheme is the one found in heterotic compactifications, where one of the factors hosts the SM sector and the other one is a hidden sector. A gaugino condensate in the latter will generate a scalar potential that will both stabilize moduli and break supersymmetry, which is then communicated to the visible sector via gravitational effects Derendinger:1985kk (123, 124). The problem with this mechanism is that the generated potential has runaway directions toward a non-interacting theory Dine:1985he (125) and so extra ingredients must be added to generate a more involved potential with actual minima. One possibility is to add background fluxes, but their quantization implies that supersymmetry is broken at a high scale. Alternatively, one may consider several gauge factors that yield a gaugino condensate, so that the generated superpotential involves a sum of competing exponentials, also known as a racetrack scenario Krasnikov:1987jj (126). This idea can be made precise in toroidal orbifold settings, since there one can use modular invariance to compute the one-loop corrections to the gauge kinetic function, and from there the exact superpotential Font:1990nt (127, 128, 129). The racetrack scenario has however the disadvantage of typically leading to vacua with negative cosmological constant.
Type II orientifold compactifications provide a different scheme. Particularly interesting are Type IIB compactifications with O3/O7-planes. Here we have two different mechanisms that generate a scalar potential, and each of them acts on a different set of light fields, with no kinetic mixing at tree-level between both sets. The first mechanism consists of adding background three-form fluxes, which fix complex structure moduli at a high scale . The second one is the presence of D-brane instantons or perturbative corrections, which stabilize Kähler moduli at a lower scale. This structure has led to several proposals to build de Sitter vacua in string theory Kachru:2003aw (117, 118), whose implementation is currently under debate. One interesting property of this scheme is that three-form fluxes can in principle break supersymmetry at a scale much lower than , and that one can compute microscopically the effect of this supersymmetry breaking on the SM sector of the compactification, which is realized on the worldvolume of D-branes Camara:2003ku (130, 131, 132). This implies that, in Type II and F-theory constructions where the MSSM or extensions are localized on a patch of the extra dimensions, one can provide a geometric description of their soft term structure.
@@
From the macroscopic viewpoint, both the heterotic and Type II models fall into the category of gravity-mediated supersymmetry breaking. This is a general 4d supergravity framework in which the MSSM soft terms are parametrized by F-terms for bulk fields, which have Planck-suppressed couplings to the SM fields Kaplunovsky:1993rd (134, 135). In the simplest cases, realized by dilaton domination in the heterotic case and by -modulus domination in intersecting D7-branes/F-theory models, all soft terms are specified by only two parameters: the universal gaugino mass and the Higgs -term. This simple pattern has the advantage that it can easily avoid unwanted flavor-changing neutral currents, because the flavor patterns in the holomorphic Yukawa sector depend on a set of moduli that decouple from the supersymmetry-breaking sector Conlon:2007dw (136). Imposing experimental constraints on these models (like compatibility with EW symmetry breaking) reduces the two parameters to one, and leads to a scenario where the eventual measurement of a single sparticle mass determines the whole spectrum Aparicio:2008wh (137). In other constructions, like -modulus domination in heterotic and D3-branes at singularities models, the leading-order soft terms vanish, and they only appear through one-loop corrections to the gauge kinetic functions and Kähler potential. Such corrections are however quite model dependent, so it is difficult to draw general lessons.
Alternatives to gravity mediated supersymmetry breaking have also been explored in string theory setups, but their full realization remains an open problem. On the one hand, one-loop anomaly mediation requires to implement the idea of sequestering in the SM sector, which has been explored in Type IIB orientifolds with the SM localized on a warped throat Kachru:2007xp (138, 139). On the other hand, gauge mediation needs to localize the SM in a small region of the extra dimensions and full stabilization of the bulk fields of the compactification at a high scale Diaconescu:2005pc (140, 141, 142). These requirements are in tension with the realizations of the SM known to date.
@@ 4.4 Dark Sectors
One might wonder whether the appearance of dark sectors is a moduli dependent statement. If the hidden gauge sector contains massless charged scalar fields, its gauge algebra can in principle be partially higgsed, thereby reducing also the amount of matter, possibly even completely. However, so-called non-Higgsable clusters Morrison:2012np (145, 146), i.e. gauge sectors with no massless charged scalars that could induce a (partial) Higgsing, seem to be another generic property of the non-perturbative landscape of string vacua. This phenomenon can be studied explicitly in F-theory compactifications: An overwhelming prevalence of non-Higgsable clusters has been observed within a large set of elliptic Calabi-Yau fourfolds Halverson:2017ffz (147, 148). The glueballs of such confining dark sectors may give rise to interesting dark matter candidates though a typical challenge in this scenario is to prevent the potential oversaturation of the relic abundance. The phenomenological aspects of this scenario have been studied, for instance, in Halverson:2018olu (149, 150) and references therein. This is just one example of dark gauge sectors with potentially interesting phenomenology (see e.g. for a different setup based on strong coupling Heckman:2011sw (151)).
The rich sequestered gauge sectors described above arise in addition to the matter sector associated with the higher-dimensional gravitational dynamics itself. Apart from the moduli and axion sector this may potentially even include the Kaluza-Klein towers of light fields such as the graviton. The idea of a light Klauza-Klein tower as the origin of dark matter, first suggested in Dienes:2011ja (152, 153), was recently investigated in Gonzalo:2022jac (154) in the context of the Swampland approach to quantum gravity. It rests on the assumption that the smallness of the observed value of the cosmological constant points to our vacuum as arising in the asymptotic regions of a moduli space with one large extra dimension Montero:2022prj (155). An alternative (and possibly dual) dark matter candidate in such scenarios are primordial black holes, see Anchordoqui:2022txe (156) and references therein.
@@ 5 THE STANDARD MODEL AND QUANTUM GRAVITY
Interestingly, a different notion of naturalness has emerged from studying the vast but finite landscape of string vacua. Stringy naturalness Douglas:2004zg (162) is a measure of the degree of tuning by the number of string theory solutions leading to a given value of an observable. Naturalness based on the statistics of string vacua has been used to assess the likelihood of the scale of supersymmetry breaking Susskind:2004uv (163, 164, 165) and in addressing other vexing hierarchy problems such as the smallness of the cosmological constant, though explicit string theory realizations of this idea remain to be found.
@@
It was in this backdrop of finding how gravity affects quantum field theory reasonings that the Swampland Program was proposed, named after the term introduced in Vafa:2005ui (166). The goal of the Swampland Program is to discriminate between EFTs that belong to the Landscape (defined to be the set of theories that can descend from a quantum theory of gravity) and those that do not. See e.g. Brennan:2017rbf (167, 168, 169, 170, 171) for recent reviews. A number of quantum gravity constraints (also known as Swampland criteria) have been put forth, mostly by finding patterns in string constructions or through thought experiments involving black holes. The motive behind finding these constraints is to discover the underlying universal principles for consistently coupling quantum field theories to gravity. These quantum gravity constraints, if proven, can have interesting phenomenological consequences. For example, milli-charged dark matter scenarios frequently considered in phenomenological model building are in tension with the absence of global symmetry in quantum gravity Shiu:2013wxa (172). The Weak Gravity Conjecture (WGC) Arkani-Hamed:2006emk (173), which quantifies how much does gravity violate global symmetries, has been used to put phenomenological constraints on axions Brown:2015iha (174, 175, 176, 177, 106) and dark photons Reece:2018zvv (178). There have been attempts to leverage the WGC in the presence of scalars to set an upper bound on the weak scale Lust:2017wrl (179, 180). A stronger versions of the WGC which presumes the absence of stable non-supersymmetric AdS vacua Ooguri:2016pdq (181) has been used to link the observed value of the cosmological constant with the neutrino masses (which for fixed Yukawa couplings, set the weak scale) Ooguri:2016pdq (181, 182, 183). It should be noted that such arguments hinge on being able to demonstrate that the AdS vacua in question are stable. Without knowledge of the full UV spectrum, we cannot exclude the possibility of lower energy solutions. In fact, the Standard Model itself already contains a source of instability. Charged matter like the electron generates a potential for the Wilson line moduli (originated from the of the Standard Model upon compactification). It was pointed out in Hamada:2017yji (183) that the potential after taking into account the Wilson line moduli has a runaway behavior. Gravitational positivity bounds Hamada:2018dde (184, 185, 186, 187) obtained through studies of scattering amplitudes have recently been used to constrain new physics models involving massive gauge bosons Aoki:2023khq (188). Since the focus of this review is to extract lessons we learned from top-down constructions rather than bottom-up Swampland considerations, we refer the readers to some recent reviews Cvetic:2022fnv (4, 189, 190) for an in-depth discussion of other phenomenological consequences of the Swampland idea.
@@
As we have seen in the previous sections, the gauge and matter content of top-down string constructions are tightly constrained. An interesting question is whether the constrained spectra are artifacts of looking under the lamppost, or there are general principles that forbid certain particles or interactions to arise in a consistent quantum theory of gravity. Recent works have gathered support for the latter. The Completeness Hypothesis Polchinski:2003bq (191) asserts that physical states with all possible gauge charges consistent with Dirac quantization are present in a consistent theory of quantum gravity. It has since been found that the Completeness Hypothesis follows from the absence of generalized non-invertible global symmetries Gaiotto:2014kfa (192, 159, 193, 194). A corollary of this hypothesis when applied to states charged under 1-form gauge fields is the existence of magnetic monopoles. When higher-form symmetries come into play, the physical states in question would include extended objects, such as strings and branes. A novel approach using brane probes to test the consistency of EFTs coupled to gravity was initiated in Kim:2019vuc (195) where it was shown that unitarity on the brane probes can rule out infinite families of anomaly free gravitational theories. Further studies have found precise match of the allowed spectra with string constructions, sometimes involving the global structure of the spacetime gauge group Cvetic:2020kuw (196, 197, 198, 199, 200, 201, 202, 203). Taken together, these findings suggest a notion of string universality Adams:2010zy (204) (or the string lamppost principle Montero:2020icj (197)) that all consistent theories of quantum gravity are realized in string theory. It is conceivable that exploring this avenue of research to gravitational theories in lower dimensions and with reduced supersymmetry may shed light on the constrained spectra observed in four-dimensional string vacua that bear significance in phenomenological contexts. Before then, six-dimensional supergravity theories and their string/F-theory realizations are an intermediate testbed for substantiating this approach. This is because this class of string constructions is fairly well understood, and known quantum consistency conditions already tightly bound the set of possible low-energy theories. See Morrison:2021wuv (205, 206) for investigations of the Completeness Hypothesis and related quantum gravity constraints in this controlled setting.
@@
We hope we have conveyed that further understanding of quantum gravity constraints, in tandem with continued advances in string compactifications, has the potential to reveal the inherent structure found in the particle spectra prevalent in string theory.
@@ 6 WHAT HAVE WE LEARNED?
Quantum gravity is essential to understand particle physics: Gravitational physics plays a more important role for particle model building than naively expected, and string theory, unlike purely field theoretic approaches to phenomenology, has all of the ingredients to help us understand this profound connection. For instance, all aspects of Abelian or discrete gauge symmetries and their associated selection rules are sensitive to the gravitational sector and cannot be studied in isolation. More generally this is a key lesson from the Swampland Program. Harvesting the constraints from quantum gravity for particle physics via string theory as a computationally well defined framework furnishes a promising research direction for modern string phenomenology.
@@ ACKNOWLEDGMENTS
We thank the editors of Annual Review of Nuclear and Particle Science for the invitation to write this short review, and Michael Peskin for useful comments. F. M. is supported through the grants CEX2020-001007-S and PID2021-123017NB-I00, funded by MCIN/AEI/10.13039/501100011033 and by ERDF A way of making Europe. G. S. is supported in part by the DOE grant DE-SC0017647. T. W. is supported in part by Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy EXC 2121 Quantum Universe 390833306 and by Deutsche Forschungsgemeinschaft through a German-Israeli Project Cooperation (DIP) grant “Holography and the Swampland”.